OpenStudy (anonymous):

The polynomial of degree 4,P(x) has a root of multiplicity 2 at x=1 and roots of multiplicity at 1 x=0 and x=-4. It goes through the point (5, 360). Find a formula for P(x).

OpenStudy (campbell_st):

well lets look st the roots so is x = 1 is a root then x - 1 = 0 then (x -1) is a factor multiplicity means that the binomial is squared x = 0 so x is a factor and lastly x = -4 so x + 4 = 0.... and (x + 4) is a factor so your polynomial is \[P(x) = a[x(x +4)(x -1)^2]\] determine the value of a, by substituting the ordered pair x = 5 P(5) = 360 and solving for a. hope this helps

OpenStudy (anonymous):

I attempted , and i have a =1/2 ?

OpenStudy (campbell_st):

yep thats the value of a... so distribute then multiply each term by 1/2 and you have the polynomial

OpenStudy (anonymous):

Okay im not sure if i wrote this correctly. But 1/2(x-1)^2(x)(x+4) ?

OpenStudy (campbell_st):

well thats basically it... so you have P(x) = 1/2[(x^2 -2x + 1)(x^2 + 4x)] hope that helps

OpenStudy (anonymous):

Yes, it does Thank YOU !